OpenStudy (anonymous):

simplify the radical expression √2/r4, all variables represent positive real numbers.

OpenStudy (anonymous):

\[\sqrt(frac{2}{r^4})\]Square root is simply exponent equal to 1/2, so the expression becomes\[(\frac{2}{r^4})^{1/2}\]\[\frac{2^{1/2}}{r^{4\times(1/2)}}\]Simplify the exponents and you'd have the final expression.

OpenStudy (anonymous):

The simplification of exponent is only required in the denominator, 1/2 in the numerator can be replaced by the square root sign again.

OpenStudy (anonymous):

okay thanks i got it.

OpenStudy (anonymous):

You are welcome :)